OpenStudy (anonymous):

Divide the following polynomial. Write your answer in order of descending powers of x. (x^3 + y^3) /(x - y)

OpenStudy (anonymous):

(x+y)(x^2+y^2−xy)/x−y is what I got

OpenStudy (anonymous):

\[x^2+x y+y^2+\frac{2 y^3}{x-y}\]

OpenStudy (anonymous):

Thank you!:)

OpenStudy (anonymous):

You are welcome.

OpenStudy (anonymous):

The website I use for math help is Mathway.com. Usually it has the correct answers but this time it didn't!

OpenStudy (anonymous):

I use the Mathematica v9 program, a very powerful tool.

OpenStudy (anonymous):



OpenStudy (anonymous):

@AJ01 What is that for?

OpenStudy (anonymous):

for the first question

OpenStudy (anonymous):

That's wrong. Thanks though.

OpenStudy (anonymous):

@robtobey What about this one? (a^4 + 4b^4) / (a^2 - 2ab + 2b^2)

OpenStudy (anonymous):

it is the same as the above

OpenStudy (anonymous):

No it's not. Your answer was wrong. Thank you though. Have a nice day.

OpenStudy (anonymous):

it is.......

OpenStudy (anonymous):

Nope. But like I said, have a nice day though!

OpenStudy (anonymous):

thanks.... i want you to know how to get the answer....

OpenStudy (anonymous):

I don't need to know lol. I just need the answer. I'll tag you in another question.

OpenStudy (anonymous):

\[a^4+4 b^4=\left(a^2-2 a b+2 b^2\right) \left(a^2+2 a b+2 b^2\right) \]